# Exploring Time Delay Interferometry Ranging as a Practical Ranging Technique in the Bayesian Framework

### 05 Mar 2024 v1

@@ Abstract
Knowledge on the inter-spacecraft optical paths (i.e. delays) is one of the key elements of time delay interferometry (TDI). Conventional method for inter-spacecraft ranging mainly relies on the pseudo-random noise (PRN) code signal modulated onto the lasers. To ensure the reliability and robustness of this ranging information, it would be highly beneficial to develop other methods which could serve as cross-validations or backups. This paper explores the practical implementation of an alternative data-driven approach - time delay interferometry ranging (TDIR) - as a ranging technique independent of the PRN signal. Distinguished from previous research, the inputs of our TDIR algorithm are interferometric measurements that have only undergone a 0.1 ms accuracy preliminary clock synchronization using the ground tracking data. This significantly relaxes the stringent requirement for clock synchronization imposed by traditional TDI procedure. In order to encode the information of time-varying light travel times and clock desynchronizations, we adopt a general polynomial parametrization to represent the delays. By framing TDIR as a Bayesian parameter estimation problem, we demonstrate our algorithm with simulated data based on the numerical orbit of Taiji. In the presence of laser frequency noise and secondary noises, the estimated median values of parameters exhibit a bias of only 5.28 ns from the “true” delays, capable of suppressing laser frequency noise to the desired level. Furthermore, we have also analysed the requirements of mitigating optical bench noise and clock noise on TDIR, and presented an illustrative example for the impact of laser locking.
@@ I Introduction
Albeit overcoming the limitations of the Earth’s scale and seismic noise, space-based GW detectors subject to laser frequency noise (LFN) which is coupled with the inequality of armlengths. LFN in the raw interferometric measurements can be 6 - 8 orders stronger than the targeted GW signals. To suppress this overwhelming noise source, the TDI technique was proposed to synthesise a virtual equal-arm interferometer in the post-processing stage [11, 12, 13]. Generally, the construction of TDI channels involves appropriately delaying and combining the readouts of interferometors, including the science (i.e. inter-spacecraft), reference and test-mass interferometors. According to the capability of LFN suppression, TDI combinations can be categorized into the first and second generations. For an ideal stationary unequal-arm constellation, perfect cancellation of LFN can be achieved by the first-generation combinations. While for LISA and Taiji whose armlengths vary at rates up to several m/s, it is necessary to employ the second-generation TDI combinations to account for both the inequality and time dependence of the armlengths [14].
@@
In addition to the interferometric measurements (also known as the beatnote signals), the inter-spacecraft optical paths (i.e. delays) constitute another crucial input for TDI. Technically, the inter-spacecraft ranging can be done by modulating the PRN code onto the lasers using approximately $1\%$ of the power [15, 16, 17, 18, 19]. The resultant PRN ranging (PRNR) may suffer from several imperfections, such as ranging noise, ambiguity, modulation noise, and on-board delays, etc. TDI places stringent requirement on the precision of delays ($\mathcal{O}(10)\sim\mathcal{O}(100)\ {\rm ns}$, depending on the TDI channels [20, 21]), so the practical application of PRNR necessitates careful processing through the combination with other ranging observables [22], based on a thorough investigation and comprehensive understanding of the technical details. Due to these technical imperfections and complexities, as well as the fundamental importance of TDI for space-based GW detection, the development of other ranging methods that are independent of PRN signal and its processing step is also desirable. Such method could serve as a risk mitigation or cross verification, and hence improve the reliability and robustness of ranging.
@@
TDIR is a data-driven approach of inter-spacecraft ranging, which does not directly measure the delays, but instead searches for the delays that minimize the residual LFN in the TDI combinations [23]. Ref. [23] proposed the basic concept of TDIR, construing it as an optimization problem aimed at minimizing the integrated power of a single TDI channel. When the delays used in TDI are close to their true values, LFN can be suppressed well below the secondary noises (i.e. the optical measurement system (OMS) noises and residual acceleration noises of the test masses). According to the principles of Bayesian statistical inference, an optimal estimator can be formulated by whitening the power of TDI channels using the noise power spectral density (PSD) of secondary noises in the Fourier domain. Researches based on this idea view TDIR as a Bayesian parameter estimation problem, which not only yields the best estimate of delays, but also outputs the confidence interval for the estimate, facilitating the assessment of ranging precision. Ref. [24] formulated the likelihood of TDIR for he first-generation Michelson channels $\{X,Y,Z\}$, ran a Markov Chain Monte Carlo (MCMC) search on the simulated data, and constrained the delays to a $\mathcal{O}(10)\ {\rm ns}$ precision. In a following up research [25], the same algorithm was extended to the second-generation TDI and time-varying delays, presuming that the delays could be parameterized by Keplerian orbital elements. No-bias estimate of the parameters was obtained on LFN-only simulations. Complementarily, a theoretical analysis on the precision of TDIR was made by Ref. [26] using the Fisher matrix formalism. In addition to these theoretical work, an experimental validation of simplified TDIR using the GRACE Follow-On satellites was described in Ref. [27].
@@
A subtle aspect about TDIR is its relationship with clock synchronization. The laser beatnotes and other measurements are recorded with respect to the clocks (such as ultra-stable oscillators, USOs) hosted by the three S/Cs, each exhibiting a unique deviation from the proper time due to instrumental imperfections including long-term deterministic drifts and random jitters. To relate the data with astrophysical events and in-orbit incidents, thereby enabling the estimation of GW source parameters and monitoring the status of detector, it is essential to adjust the timestamps of data from the three S/Cs to a common time frame (i.e. clock synchronization). In the context of TDI, ranging and synchronization are closely related for at least two reasons. Firstly, during the construction of TDI channels, the time frame of beatnote signals must be consistent with the delays applied on them. For example, if the signals are synchronized to the Solar system barycentric coordinate time (TCB) [28, 29, 30], then the delays must be the light travel time (LTT) defined as the difference of TCB times between photon reception and emission. Ref. [21, 31] predicted that TDI requires this synchronization to reach $\mathcal{O}(10)$-$\mathcal{O}(100)\ {\rm ns}$ precision. While, if the signals are the unsynchronized raw measurements, then the delays should be accordingly the PRNR, defined as the difference of onboard clock readings between photon reception and emission, which encodes the information of both LTT and clock deviations (i.e. the “TDI without synchronization” scheme, see Ref. [32] for more detailed explanations). Secondly, in the conventional data processing pipeline [33], both the nanosecond level high-precision ranging and clock synchronization rely on the PRN signal. An important aspect of PRN signal processing is to disentangle the LTTs and clock deviations using Kalman-like filters, based on the modeling of clocks, orbital motions and relativistic corrections.
@@
Most of the previous studies [23, 24, 25] illustrated TDIR on ideally synchronized data (e.g. synchronization is explicitly stated in Ref. [23], and Ref. [25] modeled the delays as LTTs, hence implicitly assuming synchronized to TCB). As is explained, these implementations of TDIR inherently relied on PRNR. Therefore, using ideally synchronized data contradicts our intention to avoid the dependence of TDIR on the PRN signal. Ref. [22] applied TDIR to unsynchronized raw measurements, while the authors only treated it as an auxiliary step in the so-called “ranging sensor fusion” processing, to correct the bias and ambiguity of PRNR. In this paper, we investigate the feasibility of TDIR on data that has been roughly synchronized to TCB with a $0.1\ {\rm ms}$ precision (still failing to meet the requirement of TDI) using the ground tracking data. As will be explained in Sec. II, in principle, TDIR also applies to completely unsynchronized data. However, for the sake of GW source parameter estimation, synchronization to TCB (or other global time frames) is inevitable. The requirement of GW analysis on clock synchronization is much less stringent than LFN suppression. For the Taiji mission, the specific requirement is under on-going research based on the scientific objectives. In the context of ranging and LFN suppression, this preliminary synchronization offers the opportunity of combining different TDI channels to achieve better constraint on the delay parameters. Accordingly, to encompass the information of both LTT and clock desynchronization into TDIR, we need to adopt a more general parametrization (other than constant delays or Keplerian orbital elements).
@@
Except for LFN, The complete TDI procedure is also known to mitigate the noises induced by the random displacements of optical benches (OBs) relative to the local inertial frames (OB noise), and by the jitters of on board clocks (clock noise). Mitigation of the former noise is accomplished through the construction of test-mass-to-test-mass interferometry (intermediate variable $\xi$), and has been included in previous studies. The latter, however, which should be dealt with an extra clock noise reduction step subsequent to the combination of TDI channels using the information of clock sidebands [34, 35], has received scant attention in the literature about TDIR. In fact, the residual clock noise in the second-generation TDI channel is still 2 - 3 orders larger than the tolerable noise level. If the clock noise reduction step is omitted, whitening the data using only the PSD of secondary noises would be incorrect. Therefore, we suggest that the practical implementation of TDIR should encompass this crucial step. Furthermore, through theoretical analysis and simulation, we investigate the requirements of suppressing OB and clock noises on the precision of TDIR.
@@ II.1 TDIR on unsynchronized data
For the applications of TDI, it is crucial that the delays used to shift the signals must be consistent with the actual delays caused by the physical propagation along the arms (i.e. the delays appearing in the equations of signals). Formally Eq. (8-11) looks similar to their “LTT versions” [18, 34, 24, 25], thus all the deduction and algorithm of TDI still applies, and the only two alterations which should be made are to change the delays from LTT to TDIR, and to interpret the beatnote signals as the unsynchronized measurements. We therefore proved that TDIR can be used to suppress LFN.
@@
It is worth noting that we have not set any restriction to the physical origin of clock deviation $\delta\hat{\tau}_{i}^{t}(\tau)$. In principle, it can be either the original (uncorrected) clock deviations, or the residual timestamp errors left by a preliminary synchronization using ground tracking data (ground-tracking synchronization hereafter), whose uncertainty is reported to be $\sim 0.1\ {\rm ms}$ [22]. Placing this preliminary synchronization step prior to TDI offers the opportunity to combine different TDI channels, so as to achieve better constraint on the delay parameters.
Besides, it is reported by Ref. [22] that orbital determination provided by ground tracking enables us to calculate LTT [40] (ground-tracking range hereafter) with an accuracy of $50\ {\rm km}$ ($\sim 0.1$ ms). This information can be used in the modeling of secondary noises, and we will demonstrate it allows to accelerate the processing of clock noise in TDIR. We set the ground-tracking ranges to have a 0.1 ms order error relative to LTTs. Our algorithm doesn’t rely on the correlations between this ranging error and the error of ground-tracking synchronization, thus they are assumed independent of each other. The differences between ground-tracking ranges and fiducial TDIR are $\delta\hat{\tau}_{i}^{t}(\tau)-\delta\hat{\tau}_{j}^{t}\left[\tau-d^{t}_{ij}(\tau)\right]$ plus the aforementioned errors of ground-tracking ranges, adding up to total differences of less than 1 ms.
To give a more intuitive explanation of the principles, beatnote signals in the above derivation are all represented in phase units. Following recent progress in the TDI algorithm [41], we perform simulation and TDI processing with data in the frequency unit. Accordingly, all the time delay operators should be replaced by the Doppler delay operators:
@@ II.2 Requirements of noise suppression on TDIR
The noise cancellation performance of TDI is usually evaluated relative to the secondary noises, mainly including the acceleration noise of test masses, corresponding to the $\delta$ terms in Eq. (8-11), and the OMS noises, corresponding to the $N^{\rm ifo}$ terms. To make an order-of-magnitude analysis, we take the $X_{2}$ channel as an example, and consider the equal-arm case. The PSD of secondary noises in $X_{2}$ reads
@@
Next, we investigate the residual noises after TDI caused by the inaccuracy of ranging. For the sake of generality, we do not set a specific laser locking scheme [42, 38] but rather assume all the six lasers are independent. Besides, to derive analytical expressions for a given level of ranging accuracy, the errors of ranging are set as constant biases dubbed $\Delta R_{ij}$. According to Eq. (12-II.1), the PSDs of the residual LFN and OB noise caused by the coupling with ranging errors in the $X_{2}$ channel are
@@
defines the average level of ranging errors. Eq. (18) is in consistency with the results of previous studies [32, 43], and one can easily verify that the $\mathcal{O}(10)\ {\rm ns}$ requirement is reasonable for the noise models in consideration. While, under the same $\Delta R$, the residual OB noise is much smaller. These expressions are verified via simulations. The algorithms of clock noise reduction varies in the literature (e.g. [34, 35, 39]), thus the residual clock noise can not be written in a unified analytical form. We adopt the treatment of Ref. [35], and explore the impact of ranging error via a clock-noise-only simulation.
Figure 2: Analytical and simulation results of the residual noises caused by ranging errors in the Michelson $X_{2}$ channel. All the ranging errors are generated from normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises. The grey, green and orange ASD plots represent the residual LFN, OB noise and clock noise, with the corresponding analytical expressions shown in black dashed line. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression.
The analytical and simulation results are visualized in FIG. 2. In our simulation, the ranging errors are generated according to normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises, which are 10 ns, 10 ns, 1 ms for the LFN, OB noise and clock noise, respectively. The grey, green and orange amplitude spectral density (ASD) plots represent the residual LFN, OB noise and clock noise, with the corresponding analytical expressions shown in black dashed line. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression. Besides the consistency between theoretical analysis and simulations, the figure also indicates that although the ground-tracking ranges deviate from the fiducial TDIR for less than 1 ms, they are already sufficient for clock noise correction. This result is beneficial for the practical implementation of TDIR. Currently the post-TDI clock noise algorithms are usually more complex than TDI itself, therefore require more computational time. To tackle this issue, we do not need to compute $X_{2,q}$ at every iteration of the MCMC run, but rather calculate it with the ground-tracking ranges in advance. Furthermore, within the regions of parameter space of our interest (tens of nanoseconds around the fiducial TDIR), OB and clock noises only make negligible contributions to the total noise budget, thus the performance of TDIR is mainly determined by LFN and the secondary noises.

### 12 Mar 2024 v3

# Exploring Time Delay Interferometry Ranging as a Practical Ranging Approach in the Bayesian Framework

@@ Abstract
Time Delay Interferometry (TDI) is an indispensable step in the whole data processing procedure of space-based gravitational wave detection, as it mitigates the overwhelming laser frequency noise, which would otherwise completely bury the gravitational wave signals. Knowledge on the inter-spacecraft optical paths (i.e. delays) is one of the key elements of TDI. Conventional method for inter-spacecraft ranging mainly relies on the pseudo-random noise (PRN) code signal modulated onto the lasers. To ensure the reliability and robustness of this ranging information, it would be highly beneficial to develop other methods which could serve as cross-validations or backups. This paper explores the practical implementation of an alternative data-driven approach - time delay interferometry ranging (TDIR) - as a ranging technique independent of the PRN signal. Distinguished from previous research, our TDIR algorithm significantly relaxes the stringent requirement for clock synchronization imposed by traditional TDI procedure. By framing TDIR as a Bayesian parameter estimation problem and employing a general polynomial parametrization, we demonstrate our algorithm with simulated data based on the numerical orbit of Taiji. In the presence of laser frequency noise and secondary noises, the estimated median values of delays are only 5.28 ns away from the ground truths, capable of suppressing laser frequency noise to the desired level. Additionally, we have also analysed the requirements of mitigating optical bench noise and clock noise on TDIR, and presented an illustrative example for the impact of laser locking.
@@ I Introduction
Albeit overcoming the limitations of the Earth’s scale and seismic noise, space-based GW detectors subject to laser frequency noise (LFN) which is coupled with the inequality of armlengths. LFN in the raw interferometric measurements can be 6 - 8 orders stronger than the targeted GW signals. To suppress this overwhelming noise source, the TDI technique was proposed to synthesise a virtual equal-arm interferometer in the post-processing stage [11, 12, 13]. Generally, the construction of TDI combinations involves appropriately delaying and combining the readouts of interferometors, including the science (i.e. inter-spacecraft), reference and test-mass interferometors. According to the capability of LFN suppression, TDI combinations can be categorized into the first and second generations. For an ideal stationary unequal-arm constellation, perfect cancellation of LFN can be achieved by the first-generation combinations. While for LISA and Taiji whose armlengths vary at rates up to several m/s, it is necessary to employ the second-generation TDI combinations to account for both the inequality and time dependence of the armlengths [14].
@@
In addition to the interferometric measurements (also known as the beatnote signals), the inter-spacecraft optical paths (i.e. delays) constitute another crucial input for TDI. Technically, the inter-spacecraft ranging can be done by modulating the PRN code onto the lasers using approximately $1\%$ of the power [15, 16, 17, 18, 19]. The resultant PRN range (PRNR) may suffer from several imperfections, such as ranging noise, ambiguity, modulation noise, and on-board delays, etc. TDI places stringent requirement on the precision of delays ($\mathcal{O}(10)\sim\mathcal{O}(100)\ {\rm ns}$, depending on the TDI channels [20, 21]), so the practical application of PRNR necessitates careful processing through the combination with other ranging observables [22], based on a thorough investigation and comprehensive understanding of the technical details. Due to these technical imperfections and complexities, as well as the fundamental importance of TDI for space-based GW detection, the development of other ranging methods that are independent of PRN signal and its processing step is also desirable. Such method could serve as a risk mitigation or cross verification, and hence improve the reliability and robustness of ranging.
@@
TDIR is a data-driven approach of inter-spacecraft ranging, which does not directly measure the delays, but instead searches for the delays that minimize the residual LFN in the TDI combinations [23]. Ref. [23] proposed the basic concept of TDIR, construing it as an optimization problem aimed at minimizing the integrated power of a single TDI channel. When the delays used in TDI are close to their true values, LFN can be suppressed well below the secondary noises (i.e. the optical measurement system (OMS) noises and residual acceleration noises of the test masses). According to the principles of Bayesian statistical inference, an optimal estimator can be formulated by whitening the power of TDI channels using the noise power spectral density (PSD) of secondary noises in the Fourier domain. Researches based on this idea view TDIR as a Bayesian parameter estimation problem, which not only yields the best estimate of delays, but also outputs the confidence interval for the estimate, facilitating the assessment of ranging precision. Ref. [24] formulated the likelihood of TDIR for he first-generation Michelson channels $\{X,Y,Z\}$, ran a Markov Chain Monte Carlo (MCMC) search on the simulated data, and constrained the delays to a $\mathcal{O}(10)\ {\rm ns}$ precision. In a following up research [25], the same algorithm was extended to the second-generation TDI and time-varying delays, presuming that the delays could be parameterized by Keplerian orbital elements. No-bias estimate of the parameters was obtained on LFN-only simulations. Complementarily, a theoretical analysis on the precision of TDIR was made by Ref. [26] using the Fisher matrix formalism. In addition to these theoretical work, the verification of a simplified TDIR algorithm using the data of GRACE Follow-On satellites was described in Ref. [27].
@@
A subtle aspect about TDIR is its relationship with clock synchronization. The laser beatnotes and other measurements are recorded with respect to the clocks (such as ultra-stable oscillators, USOs) hosted by the three S/Cs, each exhibiting a unique deviation from the proper time due to instrumental imperfections including long-term deterministic drifts and random jitters. To relate the data with astrophysical events and in-orbit incidents, thereby enabling the estimation of GW source parameters and monitoring the status of detector, it is essential to adjust the timestamps of data from the three S/Cs to a common time frame (i.e. clock synchronization). In the context of TDI, ranging and synchronization are closely related for at least two reasons. Firstly, during the construction of TDI channels, the time frame of beatnote signals must be consistent with the delays applied on them. For example, if the signals are synchronized to the Solar system barycentric coordinate time (TCB) [28, 29, 30], then the delays must be the light travel time (LTT) defined as the difference of TCB times between photon reception and emission. Ref. [21, 31] predicted that TDI requires clock synchronization to reach a $\mathcal{O}(10)$-$\mathcal{O}(100)\ {\rm ns}$ precision. While, if the signals are the unsynchronized raw measurements, then the delays should be accordingly the PRNR, defined as the difference of onboard clock readings between photon reception and emission, which encodes the information of both LTT and clock deviations (i.e. the “TDI without synchronization” scheme, see Ref. [32] for more detailed explanations). Secondly, in the conventional data processing pipeline [33], both the nanosecond level high-precision ranging and clock synchronization rely on the PRN signal. An important aspect in the processing of PRN signal is to disentangle the LTTs and clock deviations using Kalman-like filters, based on the modeling of clocks, orbital motions and relativistic corrections.
@@
Most of the previous studies [23, 24, 25] illustrated TDIR on ideally synchronized data (e.g. synchronization is explicitly stated in Ref. [23], and Ref. [25] modeled the delays as LTTs, hence implicitly assuming synchronized to TCB). As is explained, these implementations of TDIR inherently relied on PRNR. Therefore, using ideally synchronized data contradicts our intention to avoid the dependence of TDIR on the PRN signal. Ref. [22] applied TDIR to unsynchronized raw measurements, while the authors only treated it as an auxiliary step in the so-called “ranging sensor fusion” processing, to correct the bias and ambiguity of PRNR. In this paper, we investigate the feasibility of TDIR on data that has been preliminarily synchronized to TCB with a $0.1\ {\rm ms}$ precision (still failing to meet the requirement of TDI) using the ground tracking data. As will be explained in Sec. II, in principle, TDIR also applies to completely unsynchronized data. However, synchronizing to TCB (or other global time frames) is essential for the sake of GW source parameter estimation. For the Taiji mission, the precision requirement on clock synchronization imposed by astrophysical data analysis is still under on-going research guided by the scientific objectives, but it is certainly much less stringent than the that imposed by LFN suppression. Moreover, this preliminary synchronization also offers the opportunity of combining different TDI channels to achieve better constraint on the delay parameters. Accordingly, to apply TDIR on unsynchronized data, we need to adopt a more general parametrization (other than constant delays or Keplerian orbital elements).
@@
Except for LFN, The complete TDI procedure is also known to mitigate the noises induced by the random displacements of optical benches (OBs) relative to the local inertial frames (i.e. OB noises), as well as the noises originating from the jitters of on board clocks (i.e. clock noises). Mitigation of the former noise is accomplished through the construction of test-mass-to-test-mass interferometry (intermediate variable $\xi$), and has been included in previous studies. The latter, however, which should be dealt with an extra clock noise reduction step subsequent to the combination of TDI channels using the information of clock sidebands [34, 35], has received scant attention in the literature about TDIR. In fact, the residual clock noise in the second-generation TDI channel is still 2 - 3 orders larger than the tolerable noise level. If the clock noise reduction step is omitted, whitening the data using only the PSD of secondary noises would be incorrect. Therefore, we suggest that the practical implementation of TDIR should encompass this crucial step. Furthermore, through theoretical analysis and simulation, we investigate the requirements of suppressing OB and clock noises on the precision of TDIR.
@@ II.1 TDIR on unsynchronized data
For the applications of TDI, it is crucial that the delays used to shift the signals must be consistent with the delays appearing in the equations of signals. Formally Eq. (8-11) looks similar to their “LTT versions” [18, 34, 24, 25], thus all the deduction and algorithm of TDI still applies, and the only two alterations which should be made are to change the delays from LTT to TDIR, and to interpret the beatnote signals as the unsynchronized measurements. We therefore proved that TDIR can be used to suppress LFN on unsynchronized data.
@@
It is worth noting that we have not set any restriction to the physical origin of clock deviation $\delta\hat{\tau}_{i}^{t}(\tau)$. In principle, it can be either the original (uncorrected) clock deviations, or the residual timestamp errors left by a preliminary synchronization using ground-tracking data (ground-tracking synchronization hereafter), whose uncertainty is reported to be $\sim 0.1\ {\rm ms}$ [22]. In this paper we explore the application of TDIR in the second scenario. Placing this preliminary synchronization step prior to TDI offers the opportunity to combine different TDI channels, so as to achieve better constraint on the delay parameters.
Besides, it is reported by Ref. [22] that the orbital determination provided by ground tracking enables us to calculate LTTs [40] (ground-tracking ranges hereafter) with an uncertainty of $50\ {\rm km}$ ($\sim 0.1$ ms). This information can be used to model the secondary noises in the calculation of TDIR likelihood (see Sec. III), and we will demonstrate that it allows to accelerate the processing of clock noise in TDIR. The differences between ground-tracking ranges and fiducial TDIR are $\delta\hat{\tau}_{i}^{t}(\tau)-\delta\hat{\tau}_{j}^{t}\left[\tau-d^{t}_{ij}(\tau)\right]$ plus the aforementioned uncertainties of ground-tracking ranges, adding up to total differences of less than 1 ms.
To give a more intuitive explanation of the principles, beatnote signals in the above derivation are all represented in phase units. Following recent progress in the TDI algorithm [41], we perform TDI processing with simulated data in the frequency unit. Accordingly, all the time delay operators should be replaced by the Doppler delay operators:
@@ II.2 Requirements of noise suppression on TDIR
The noise cancellation performance of TDI is usually evaluated relative to the secondary noises, mainly including the acceleration noise of test masses, corresponding to the $\delta$ terms in Eq. (8-11), and the OMS noises, corresponding to the $N^{\rm ifo}$ terms. To make an order-of-magnitude analysis, we take the $X_{2}$ channel as an example, and consider the equal-arm case. The PSD of secondary noises reads
@@
Next, we investigate the residual LRN, OB noise and clock noise caused by the ranging errors. For generality, we do not set a specific laser locking scheme [42, 38] but rather assume all the six lasers are independent. Besides, to derive analytical expressions for a given level of ranging accuracy, the errors of ranging are set as constant biases dubbed $\Delta R_{ij}$. According to Eq. (12-II.1), the PSDs of the residual LFN and OB noise in the $X_{2}$ channel are
@@
defines the average ranging error. Eq. (18) is in consistency with the results of previous studies [32, 43], and one can easily verify that a $\mathcal{O}(10)\ {\rm ns}$ requirement is reasonable for the noise models in consideration. While, under the same $\Delta R$, the residual OB noise is much smaller. The algorithms of clock noise reduction varies in the literature (e.g. [34, 35, 39]), thus the residual clock noise can not be written in a unified analytical form. We adopt the treatment of Ref. [35], and explore the impact of ranging error via a clock-noise-only simulation.
Figure 2: Analytical and simulation results of the residual noises caused by ranging errors in the Michelson $X_{2}$ channel. All the ranging errors are generated from normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises. The grey, green and orange ASD plots represent the simulated residual LFN, OB noise and clock noise, with the corresponding theoretical results shown in black dashed curves. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression.
The analytical and simulation results are visualized in FIG. 2. In our simulation, the ranging errors are generated according to normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises, which are 10 ns, 10 ns, 1 ms for the LFN, OB noise and clock noise, respectively. The grey, green and orange amplitude spectral density (ASD) plots represent the simulated residual LFN, OB noise and clock noise, with the corresponding theoretical results shown in black dashed curves. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression. Besides the consistency between theoretical analysis and simulations, the figure also indicates that although the ground-tracking ranges deviate from the fiducial TDIR for less than 1 ms, they are already sufficient for suppressing the clock noises. This result is beneficial for the practical implementation of TDIR. Currently the post-TDI clock noise algorithms are usually more complex than TDI itself, therefore require more computational time. To tackle this issue, we do not need to compute $X_{2,q}$ at every iteration of the MCMC run, but rather calculate it with the ground-tracking ranges in advance. Furthermore, within the regions of parameter space of our interest (tens of nanoseconds around the fiducial TDIR), OB and clock noises only make negligible contributions to the total noise budget, thus the performance of TDIR is mainly determined by LFN and the secondary noises.